OpenStudy (anonymous):

integral of sec^2((pi/16)x)-1dx

zepdrix (zepdrix):

Hi, welcome to OpenStudy.\[\Large\rm \int\limits \sec^2\left(\frac{\pi}{16}x\right)-1~~dx\]So we need to integrate this?

zepdrix (zepdrix):

Do you remember this integral? \(\Large\rm \int sec^2x~dx\) Or maybe I should ask... do you remember this derivative? \(\Large\rm \dfrac{d}{dx}\tan x=\sec^2x\)

OpenStudy (anonymous):

yes i remember that but doesn't the -1 effect it?

OpenStudy (anonymous):

\[\int sec^2(something)-1 dx=\int sec^2(something)dx -\int 1dx\]